Please help me solve this function: Determine rather the function is even,od,or neither: f(x)= x^5 - x
Answer by lwsshak3(11628)

Please help me solve this function: Determine rather the function is even,odd,or neither: f(x)= x^5-x
If f(-x)=f(x) function is even
If f(-x)=-f(x) function is odd
For given function:
f(x)= x^5-x
f(x)=x(x^4-1)
f(-x)=-x(x^4-1)
f(-x)=-f(x)
Therefore, function is odd